OpenStudy (amistre64):

How many vectors are in the in the \(span\left{\begin{pmatrix}7&-4&-6\end{pmatrix},\begin{pmatrix}-8&5&7\end{pmatrix},\begin{pmatrix}7&0&-5\end{pmatrix}\right}\) there has GOT to be an easier way to code up column vectors .....

OpenStudy (amistre64):

\[ span\left{ \begin{pmatrix}7&-4&-6 \end{pmatrix}, \begin{pmatrix}-8&5&7 \end{pmatrix}, \begin{pmatrix}7&0&-5 \end{pmatrix}\right}\]

OpenStudy (amistre64):

yeah, this thing doesnt code well ....

OpenStudy (amistre64):

7 -8 7 -4 5 0 -6 7 -5 are the vectors in the span

OpenStudy (experimentx):

how do you describe the system ... i would say 3 vectors with 3 components

OpenStudy (amistre64):

I thought 3 vectors as well; but apparently my error is in interpreting the terminology ...

OpenStudy (amistre64):

how many vectors are in the set {v1,v2,v3}? 3 how many vectors are in the span{v1,v2,v3}? infinite

OpenStudy (kinggeorge):

unless your vector field only has a finite number of elements. Then it's not infinite. However, assuming you're in \(\mathbb{R}^3\), you're right.

OpenStudy (amistre64):

If the question had asked: How many vectors can be made from span{v1,v2,v3} id have understood it better :)

OpenStudy (kinggeorge):

On one of my tests a few weeks ago, not a single person in the class used the "correct" method to find the answer to a problem. The teacher ended up admitting his wording was quite terrible.